OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1.

OpenStudy (akashdeepdeb):

The slope-intercet equation = y = mx + b Where b = y-intercept and m = slope! Substitute the values in the line equation for the answer! :)

OpenStudy (anonymous):

? huh

OpenStudy (akashdeepdeb):

y = mx + b Here b = -1 m = -5 So line equation = y = -5x -1 Or 5x + y + 1 = 0 Getting it? :)

OpenStudy (anonymous):

so what would i do next in order to get the answer and write the equation in the box x

OpenStudy (anonymous):

please just give me the answer and equation

OpenStudy (akashdeepdeb):

I already did! :) But didn't you understand it? :D

OpenStudy (anonymous):

just write the answer and equation in the thing please

OpenStudy (akashdeepdeb):

5x + y + 1 = 0 :)

OpenStudy (texaschic101):

y = -5x - 1 is slope intercept form